The formula for calculating the sum of all natural integers from 1 to n is well-known:

Sn = 1 + 2 + 3 + ... + n =

n

2 + n

2

Similary, we know about the formula for calculating the sum of the first n squares:

Qn = 1 · 1 + 2 · 2 + 3 · 3 + ... + n · n =

n

3

3

+

n

2

2

+

n

6

Now, we reduce one of the two multipliers of each product by one to get the following sum:

Mn = 0 · 1 + 1 · 2 + 2 · 3 + 3 · 4 + ... + (n − 1) · n

Find an explicit formula for calculating the sum Mn.


1
Expert's answer
2021-11-26T00:04:43-0500
"\displaystyle\sum_{i=1}^ni=\dfrac{n(n+1)}{2}"

"\displaystyle\sum_{i=1}^ni^2=\dfrac{n(n+1)(2n+1)}{6}"

"\displaystyle\sum_{i=1}^n(i-1)i=\displaystyle\sum_{i=1}^n(i^2-i)=\displaystyle\sum_{i=1}^ni^2-\displaystyle\sum_{i=1}^ni"

"=\dfrac{n(n+1)(2n+1)}{6}-\dfrac{n(n+1)}{2}"

"=\dfrac{n(n+1)(2n+1-3)}{6}"

"=\dfrac{(n-1)n(n+1)}{3}"

"M_n=0 \u00b7 1 + 1 \u00b7 2 + 2 \u00b7 3 + 3 \u00b7 4 + ... + (n \u2212 1) \u00b7 n"

"=\displaystyle\sum_{i=1}^n(i-1)i=\dfrac{(n-1)n(n+1)}{3}"